OpenStudy (anonymous):

what is the standard form of the equation that has a slope of 2/3 and passes through the point (1,0) ?

OpenStudy (anonymous):

use y=mx+b to solve 0=2/3(1)+b b=-2/3 y=2/3x-2/3

OpenStudy (anonymous):

y=(2/3)x-(2/3)

OpenStudy (anonymous):

ahh but standard form would be -2/3x+y=-2/3

OpenStudy (anonymous):

Form questions are questions that test your vocabulary as much as your math. GENERAL FORM: Ax + By + C = 0 STANDARD FORM: Ax + By = C POINT SLOPE FORM first y minus second y = (first y minus second y / first x minus second x)(first x minus second x). BY THE WAY: Remember m is (first y minus second y / first x minus second x) SLOPE INTERCEPT FORM: y = mx + b